OpenStudy (anonymous):

Find the 7th and nth term of the sequence 1,-x, x^2 i think i need to use geometric sequence here but i don't know where to start please help..

OpenStudy (anonymous):

start with -1 and an x ....

OpenStudy (anonymous):

(-1)^0 = 1 (-1)^1 = -1 (-1)^2 = 1 (-1)^3 = -1 (-1)^4 = 1 ... that gets you alternating signs, and the rest is just xs

OpenStudy (anonymous):

(-x)^n seems suitable to me